OpenStudy (anonymous):

what is the integration of f(x)= 7 integral of e^x ?

OpenStudy (anonymous):

7 integral of e^x

OpenStudy (anonymous):

i mean this

OpenStudy (anonymous):

The answer is 7 integral of e^x

OpenStudy (anonymous):

can you tell me the reason please?

OpenStudy (anonymous):

What's the integral of e^x?

OpenStudy (cwrw238):

e is a special number ( appr ox 2.718) the graph of y = e^x has the property that the value of y at all points on the graph are equal to the slope atthat point thats is d(e^x) / dx = e^x so the integral of e^x is also e^x

OpenStudy (anonymous):

$$f(x)=7\int e^x\,\mathrm{d}x=7e^x+C_1\\\int f(x)\,\mathrm{d}x=\int\left(7e^x+C_1\right)\,\mathrm{d}x=7e^x+C_1x+C_2$$